# yijia#1 [Yijia: #1] Large Language Models Can Be Good Privacy Protection Learners

### v1

@@ Abstract
The proliferation of Large Language Models (LLMs) has driven considerable interest in fine-tuning them with domain-specific data to create specialized language models. Nevertheless, such domain-specific fine-tuning data often contains sensitive personally identifiable information (PII). Direct fine-tuning LLMs on this data without privacy protection poses a risk of leakage. To address this challenge, we introduce Privacy Protection Language Models (PPLM), a novel paradigm for fine-tuning LLMs that effectively injects domain-specific knowledge while safeguarding data privacy. Our work offers a theoretical analysis for model design and delves into various techniques such as corpus curation, penalty-based unlikelihood in training loss, and instruction-based tuning, etc. Extensive experiments across diverse datasets and scenarios demonstrate the effectiveness of our approaches. In particular, instruction tuning with both positive and negative examples, stands out as a promising method, effectively protecting private data while enhancing the model’s knowledge. Our work underscores the potential for Large Language Models as robust privacy protection learners.
@@ 1 Introduction
Background. Large Language Models (LLMs) have demonstrated remarkable linguistic comprehension and generation capability (Bang et al., 2023; Wang et al., 2023a). Meanwhile, when directly applied to specialized industries, they encounter challenges such as hallucination (Bang et al., 2023; Chan et al., 2023), insufficient domain expertise (Singhal et al., 2023b), and failing to incorporate the latest domain knowledge in ever-evolving industry scenarios Kasneci et al. (2023). The introduction of open-source general-purpose LLMs such as LLaMA (Touvron et al., 2023) and RWKV (Peng et al., 2023) have provided a promising solution. Researchers would fine-tune specialized LLMs based on powerful general-purpose LLMs using high-quality, domain-specific knowledge to ensure both commonsense reasoning and comprehensive knowledge coverage(Hoffmann et al., 2022a; Villalobos et al., 2022; Hoffmann et al., 2022b). Such examples include BloombergGPT (Wu et al., 2023) and Med-PaLM (Singhal et al., 2023a), for financial and medical applications, respectively. However, these fine-tuning datasets usually contain sensitive information, such as personally identifiable information (PII) (Carlini et al., 2020; Lin et al., 2021; Gehman et al., 2020). When applied to downstream tasks, sensitive information in the training data, such as social security numbers or patient names, can be exposed by the LLMs upon text generation, a phenomenon known as the memorization effect (Yu et al., 2023b; Kenton & Toutanova, 2019; Meng et al., 2023), leading to identity theft and financial losses (Coavoux et al., 2018; Yu et al., 2023a)
@@
Challenges. In this work, we aim to tackle the challenging task of efficient LLM fine-tuning for enhanced contextual privacy, a critical yet under-explored setting where the sensitivity of a piece of information is contingent upon the context. For example, statements such as “Bill Gates founded Microsoft” and “Alan Mathison Turing was an English mathematician and computer scientist” are generally not considered violations of privacy, since they are presented as common knowledge. In contrast, statements like “Alan Gates visited the X hospital for a certain disease Y” pose privacy concerns as they reveal details about individuals’ daily activities and health status in a particular context. Directly applying techniques like Named Entity Recognition (NER) can lead to inaccurate identification of PII, whereas merely deleting or masking PII tokens in the training data would result in a substantial information loss and compromise the performance on downstream tasks — a conundrum known as the privacy-utility trade-off as theoretically discussed in Sec. 4.1. An alternative approach, reinforcement learning from human feedback (RLHF), involves additional model fine-tuning guided by human feedback (Ouyang et al., 2022) so that the model tends towards concealing sensitive PII (like “red-teaming”). For example, it learns to prioritize outputs that protect sensitive PII over those that leak PII. Nonetheless, RLHF is data-intensive, potentially costly in computation, and can pose stability challenges (Ziegler et al., 2020; Wang et al., 2023b).
@@
This Work. To address these challenges, this paper introduces effective and efficient methodologies for fine-tuning LLMs to incorporate domain knowledge while ensuring privacy protection. We propose and rigorously examine a diverse suite of strategies from corpus curation, introduction of penalty-based unlikelihood into the training loss, instruction-based tuning, a PII contextual classifier, and direct preference optimization (DPO), etc. The ultimate objective is to cultivate a model that excels at acquiring information while demonstrating the ability to distinguish between information that can be openly shared and that demands strict confidentiality. Our experimental findings suggest that instruction tuning with positive and negative examples can offer promising avenues. It not only effectively shields private data but also enables the model to assimilate knowledge from the corpus. This implies that LLMs can be good privacy protection learners, without the need for balancing a privacy-utility trade-off.
To sum up, our contributions are as follows:
@@
Novel Methodology. For the first time, we explicitly address the challenging problem of building Privacy Protection Language Models (PPLM), a novel paradigm in fine-tuning language models that emphasizes privacy protection. To achieve this, we systematically lay out and empirically test a comprehensive spectrum of strategies.
@@
Theoretical Guidance. We provide a theoretical analysis of our proposed methodologies. This analysis illuminates the pathway to designing robust tuning methods, ensuring the resultant language model can both protect private data and assimilate vast knowledge from the fine-tuning corpus.
@@
Extensive Experiments. We comprehensively evaluate our methods on three real-world datasets. These experiments demonstrated the efficacy of our fine-tuning method to inject domain knowledge and safeguard private personal information (PII). The outcomes show that our technique performs significantly better than the straightforward baselines.
Our code and data are available at Anonymous GitHub111https://anonymous.4open.science/r/PPLM. We will make all code and the proposed datasets publicly available upon the acceptance of this work.
2 Related Works
Large Language Models and Privacy. In the rapidly advancing domain of artificial intelligence and natural language processing, Large Language Models (LLMs) have emerged as a significant force, demonstrating unprecedented capabilities in generating coherent and contextually accurate text. Many LLMs released, such as GPT-3.5/4(OpenAI, 2023), Bard (Google, 2023), LLaMA (Touvron et al., 2023), and ChatGLM (Du et al., 2022). Large language models also have been applied to finance (Yang et al., 2023; Wu et al., 2023), biology (Fang et al., 2022; Rives et al., 2019) and clinical area (Singhal et al., 2023a; Yang et al., 2022). Nevertheless, it is inevitable that the application of LLM poses risks to personal information. To tackle the privacy leakage issues of LLMs, efforts have been made from three aspects (Li et al., 2023; Zhang et al., 2023): 1). curation of the pretraining corpus; 2). conditional pretraining of LLMs, and 3). alignment of LLMs after pretraining, which are discussed in the following. Note that the task is different from the Differential Privacy (DP) LLMs (Yu et al., 2022; Shi et al., 2021; Anil et al., 2022; Li et al., 2022; Liu et al., 2023) which is targeting to protect general differential privacy Zhao et al. (2022), our focus is the contextual privacy protection finetuning for knowledge injection, in which the PII to be protected in the tuning text data is annotated by users.
@@ 2 Related Works
For the pretraining corpus, manually detecting and filtering out/revising the corpus can offer high-quality corpus, which is ideal for training privacy-preserving LLMs (Hoffmann et al., 2022b; Villalobos et al., 2022). Nevertheless, it is infeasible to process billions of tokens manually in practice. Another solution is using automated tools to filter out all sensitive content (e.g. names, addresses, phone numbers) from the pretraining corpus. Automated filters make it possible to go over pretraining datasets, however, simply removing or masking the PII tokens can cause information loss or inconsistency in the corpus (Welbl et al., 2021). Though filters can ‘clean’ datasets, they reduce the diversity in the corpus, which further negatively impacts the robustness of LLMs (Hendrycks et al., 2019). Another solution is adding content filters on top of the existing LMs to control the content generation process (Xu et al., 2020). Even so, carefully designed cases (e.g. prompts) can still trigger some undesired behaviors of large LMs (Gehman et al., 2020; Ziegler et al., 2022). However, directly removing PII from the training corpus poses a dilemma. While it ensures the elimination of sensitive data, it also potentially weakens the LLMs by stripping them of crucial knowledge. The mere act of omitting data can inadvertently hamper the model’s capacity to process and understand certain contexts. Context-awareness is fundamental when considering privacy protection and what data to shield. Direct filtering without understanding the context can thus be misleading and ineffective.
Pretraining with preferences. Another solution is to maintain the content, but use redesigned loss/conditional tags to control the information injected into the LLMs. Pretraining language models with human preferences indicate to the model what information should be learned and what should be avoided, rather than forcing them to forget the learned content. Pretraining with conditional human preference scores can offer a Paterto-optimal and simple approach to reduce the undesirable content by up to an order of magnitude. Korbak et al. (2023) compared with the classical pretraining approach. While pretraining LLMs conditioned under annotation scores can offer better performance in the human preferences aspect. Since human preferences are injected into the models during the pretraining stage, the models are biased toward those preferences once they are trained. With the expanding size of LLMs, they become increasingly resistant to forgetting their training data (Carlini et al., 2022; Vu et al., 2022; Ramasesh et al., 2022; Korbak et al., 2023). In other words, pretraining large language models conditioned under preference score sacrifices some flexibility. Still, it is undeniable that it can provide much better alignment with human preferences compared with the classical pretraining schema.
LLMs adaptation. To strike a balance between performance and flexibility, pretraining large LMs without constraints and then adjusting them to align with human preferences is a widely adopted approach for now. One approach is supervising fine-tuning. The pre-trained LMs are tuned on curated datasets in a supervised manner (Solaiman & Dennison, 2021; Zhou et al., 2023; Wan et al., 2023; Jin et al., 2023). Another approach is reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022; Menick et al., 2022; Chen et al., 2023). RLHF gathers data with feedback/preference labels, trains a reward model, and then finetunes the LM with reinforcement learning.
@@ 3 Problem Statement
Problem formulation. In the context of language models, a fine-tuning dataset is a collection of natural language sequences . Each sequence is denoted as , where represents a token. For privacy protection, the users annotate each sequence in the corpus by a binary sequence denoted as , where denotes the token is private tokens (e.g., PII) need to be protected in the context, and otherwise. Here, the contextual privacy posits that the sensitivity of a piece of information is not solely intrinsic to the information itself, but is also influenced by its surrounding context. To illustrate, “Alan Gates visited Crescent Vale Medical Center for Hemophilia treatment” is considered more indicative than “Alan Gates visited Crescent Vale.” The former provides a clearer insight into an individual’s health when the name “Alan Gates” is paired with the medical condition and the specific medical center. Important notations used in the paper are included in Table 4 in the Appendix.
@@ 4 Methodology
Our methodology adopts a two-pronged approach: 1) corpus curation (i.e. filtering), where sensitive data such as personally identifiable information (PII) is removed from the corpus; and 2) supervised fine-tuning. We commence with a theoretical analysis of the information loss incurred by the corpus curation strategy, which provides guidelines for method development. Then, we propose five novel strategies for privacy protection when fine-tuning large language models.
@@ 4.1 Theoretical Analysis on the Information Loss During Corpus Curation
The alternative approach is to mask the text sequence by substituting the word with a special token wherever , then train the model to directly predict the new sequence . Here, denotes a PII token associated with sensitive information like names, organizations, addresses, and website URLs. Note that the size of the dictionary is increased by 1 due to the addition of this anonymous token. The masking procedure above is a one-way mapping from to . We denote this masking mapping as and . The revised maximum log-likelihood estimator is as follows:
@@
This implies that the right-hand side of Eq. 4.1 is larger than the right-hand side of Eq. 4.1. Therefore, directly learning offers richer information. Minimizing Eq. 4.1 ensures the value in Eq. 4.1 remains small, whereas the reverse does not hold. Overall, instructing the model with the “correct” information is more effective and informative than imposing constraints to selectively hide or forget previously acquired knowledge, such as intentionally removing or masking PIIs in the training text.
@@ 4.2.2 Penalty-Based Loss
where and are the penalty terms for generating unigrams and bigrams associated with PII. is the likelihood of generating the token associated with PII at position . is the set of n-grams associated with PII. To construct , we extract all PII-associated n-grams from the training set using scrubadub222https://github.com/LeapBeyond/scrubadub. The cumulative loss is then calculated as:
@@ 4.2.5 DPO
Compared to RLHF, DPO (Rafailov et al., 2023) eliminates the need to train a reward model, and optimizes the same objective as in RLHF with a single stage of policy training using the following objective:
@@ 5.1 Datasets
We adopt three datasets from the biomedical field: pii-medical_flashcards, pii-wikidoc and pii-wikidoc_patient_information as summarized in Table 5 in the Appendix A.1. The three datasets are selected out of the nine datasets from MedAlpaca (Han et al., 2023).
@@ 5.2 Experimental Setup
For each method, we adapt the Alpaca-style tuning pipeline of LLaMA2 (Touvron et al., 2023), from llama-receipe333https://github.com/facebookresearch/llama-recipes/. In our experiments, all the methods share the same training settings. The number of training epochs is set to 5 and the batch size is set to 64. For a fair comparison, we adopt the same backbone LLM LLaMA2 for fine-tuning. More implementation details are included in the Appendix C.

### v2

# Large Language Models Can Be Contextual Privacy Protection Learners

@@ Abstract
The proliferation of Large Language Models (LLMs) has driven considerable interest in fine-tuning them with domain-specific data to create specialized language models. Nevertheless, such domain-specific fine-tuning data often contains contextually sensitive personally identifiable information (PII). Direct fine-tuning LLMs on this data without privacy protection poses a risk of data leakage of sensitive PII during inference time. To address this challenge, we introduce Contextual Privacy Protection Language Models (CPPLM), a novel paradigm for fine-tuning LLMs that effectively injects domain-specific knowledge while safeguarding inference-time data privacy. Our work offers a theoretical analysis for model design and benchmarks various techniques such as corpus curation, penalty-based unlikelihood in training loss, instruction-based tuning, etc. Extensive experiments across diverse datasets and scenarios demonstrate the effectiveness of our approaches. In particular, instruction tuning with both positive and negative examples, stands out as a promising method, effectively protecting private data while enhancing the model’s knowledge. Our work underscores the potential for Large Language Models as robust contextual privacy protection learners. The complete code and data for the work can be found at https://github.com/Yijia-Xiao/PPLM.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable linguistic comprehension and generation capability (Bang et al., 2023; Wang et al., 2023a). Meanwhile, when directly applied to specialized industries, they encounter challenges such as hallucination (Chan et al., 2023; Deng et al., 2024; Jin et al., 2024a), insufficient domain expertise (Singhal et al., 2023b), and failing to incorporate the latest domain knowledge in ever-evolving industry scenarios (Kasneci et al., 2023). The introduction of open-source general-purpose LLMs such as LLaMA (Touvron et al., 2023) and RWKV (Peng et al., 2023) have provided a promising solution. Researchers would fine-tune specialized LLMs based on powerful general-purpose LLMs using high-quality, domain-specific knowledge to ensure both commonsense reasoning and comprehensive knowledge coverage (Hoffmann et al., 2022a, b; Villalobos et al., 2022; Yang et al., 2024). Such examples include BloombergGPT (Wu et al., 2023) and Med-PaLM (Singhal et al., 2023a), for financial and medical applications, respectively. However, these fine-tuning datasets usually contain sensitive information, such as personally identifiable information (PII) (Carlini et al., 2020; Lin et al., 2021; Gehman et al., 2020). When applied to downstream tasks, sensitive information in the training data, such as social security numbers or patient names, can be exposed by the LLMs upon text generation, a phenomenon known as the memorization effect (Yu et al., 2023b; Kenton and Toutanova, 2019; Meng et al., 2023) or inference-time privacy threat (Mireshghallah et al., 2024), leading to identity theft and financial losses (Coavoux et al., 2018; Yu et al., 2023a).
@@
Challenges. In this work, we aim to tackle the challenging task of efficient LLM fine-tuning for enhanced contextual privacy (Nissenbaum, 2004; Mireshghallah et al., 2024), a critical yet under-explored setting where the sensitivity of a piece of information is contingent upon the context. For example, statements such as “Bill Gates founded Microsoft” and “Alan Mathison Turing was an English mathematician and computer scientist” are generally not considered violations of privacy, since they are presented as common knowledge. In contrast, statements like “Alan Gates visited the X hospital for a certain disease Y” pose privacy concerns as they reveal details about individuals’ daily activities and health status in a particular context. Directly applying techniques like Named Entity Recognition (NER) can lead to inaccurate identification of PII, whereas merely deleting or masking PII tokens in the training data would result in a substantial information loss and compromise the performance on downstream tasks — a conundrum known as the privacy-utility trade-off as theoretically discussed in Sec. 4.1. An alternative approach, reinforcement learning from human feedback (RLHF), involves additional model fine-tuning guided by human feedback (Ouyang et al., 2022) so that the model tends towards concealing sensitive PII (like “red-teaming”). For example, it learns to prioritize outputs that protect sensitive PII over those that leak PII. Nonetheless, RLHF is data-intensive, potentially costly in computation, and can pose stability challenges (Ziegler et al., 2020; Wang et al., 2023b).
@@
Our Work. To address these challenges, this paper introduces effective and efficient methodologies for fine-tuning LLMs to incorporate domain knowledge while ensuring privacy protection. We propose and rigorously examine a diverse suite of strategies from corpus curation, introduction of penalty-based unlikelihood into the training loss, instruction-based tuning, a PII contextual classifier, and direct preference optimization (DPO), etc. The ultimate objective is to cultivate a model that excels at acquiring information while demonstrating the ability to distinguish between information that can be openly shared and that demands strict confidentiality. Our experimental findings suggest that instruction tuning with positive and negative examples can offer promising avenues. It not only effectively shields private data but also enables the model to assimilate knowledge from the corpus. This implies that LLMs can be good contextual privacy protection learners, without the need for balancing a privacy-utility trade-off. To sum up, our contributions are as follows.
@@
Novel Methodology. For the first time, we explicitly address the challenging problem of building Contextual Privacy Protection Language Models (CPPLM), a novel paradigm in fine-tuning language models that emphasizes privacy protection of contextual PII. To achieve this, we systematically lay out and empirically test a comprehensive spectrum of strategies.
@@
Theoretical Guidance. We provide a theoretical analysis of our proposed methodologies. This analysis illuminates the pathway to designing robust tuning methods, ensuring the resultant language model can both protect private data and assimilate vast knowledge from fine-tuning corpus.
@@
Comprehensive Evaluation. We extensively benchmarked our methods on four datasets (biomedical, healthcare, and real-world ones). These experiments demonstrated the efficacy of our fine-tuning method to inject domain knowledge and safeguard private personal information (PII). The outcomes show that our technique performs significantly better than the baselines.
2 Related Work
Large Language Models and Privacy. In the rapidly advancing domain of artificial intelligence and natural language processing, LLMs such as GPT-3.5/4 (OpenAI, 2023), Bard (Google, 2023), LLaMA (Touvron et al., 2023), and ChatGLM (Du et al., 2022) have demonstrated unprecedented capabilities in following instructions Lou et al. (2024, 2023) and generating coherent, contextually accurate text Wang et al. (2024); Xiong et al. (2024); Jiang et al. (2024a, b); Hong et al. (2024). However, this widespread application raises significant privacy concerns, particularly regarding personal information protection. Addressing the privacy challenges posed by LLMs, researchers have focused on three primary strategies: (Li et al., 2023; Zhang et al., 2023; Kim et al., 2023; Lukas et al., 2023): 1) curation of the pretraining corpus, 2) conditional large language model (LLM) pretraining, and 3) post-training alignment. Our research focuses on enhancing privacy protection in LLMs through fine-tuning methods that enable knowledge injection to safeguard Personally Identifiable Information (PII) (Lukas et al., 2023), as designated by users. This contrasts with Differential Privacy (DP), which protects against the leakage of entire records at the cost of reduced data utility Yu et al. (2022b). Our method emphasizes targeted PII protection, a crucial aspect in contexts where knowledge integration is the key to preserving privacy without compromising data utility (Shi et al., 2021; Anil et al., 2022; Li et al., 2022a; Liu et al., 2024; Zhao et al., 2022; Li et al., 2022b; Yu et al., 2022a).
For the fine-tuning of LLMs, the decline in utility is inversely linked to the privacy budget allocated for safeguarding the entire training document, as it determines the extent of noise introduced Lukas et al. (2023). Our emphasis lies in specifically safeguarding the contextual PII tokens. Since PIIs are contextual (Mireshghallah et al., 2024; Nissenbaum, 2004), our approach tunes LLMs with contrastive examples designated by users can accommodate the customized privacy preferences.
@@ 2 Related Work
For the pretraining corpus, manually detecting and filtering out/revising the corpus can offer high-quality corpus, which is ideal for training privacy-preserving LLMs (Hoffmann et al., 2022b; Villalobos et al., 2022; Lukas et al., 2023). Nevertheless, it is infeasible to process billions of tokens manually in practice. Another solution is using automated tools to filter out all sensitive content (e.g. names, addresses, phone numbers) from the pretraining corpus. Automated filters make it possible to go over pretraining datasets. However, simply removing or masking the PII tokens (i.e., PII scrubbing) can cause information loss or inconsistency in the corpus (Welbl et al., 2021). Though filters can ‘clean’ datasets, they reduce the diversity in the corpus, which further negatively impacts the robustness of LLMs (Hendrycks et al., 2019). Another solution is adding content filters on top of the existing LMs to control the content generation process (Xu et al., 2020). Even so, carefully designed cases (e.g. prompts) can still trigger some undesired behaviors of large LMs (Gehman et al., 2020; Ziegler et al., 2022). However, directly removing PII from the training corpus poses a dilemma. While it ensures the elimination of sensitive data, it also potentially weakens the LLMs by stripping them of crucial knowledge. The mere act of omitting data can inadvertently hamper the model’s capacity to process and understand certain contexts. Context-awareness is fundamental when considering privacy protection and what data to shield.
LLMs Adaptation. To strike a balance between performance and flexibility, pretraining large LMs without constraints and then adjusting them to align with human preferences is a widely adopted approach for now. One approach is supervising fine-tuning. The pre-trained LMs are tuned on curated datasets in a supervised manner (Solaiman and Dennison, 2021; Zhou et al., 2023; Wan et al., 2023; Jin et al., 2024b; Xiao et al., 2024; Lu et al., 2024). Another approach is reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022; Menick et al., 2022; Zhang et al., 2024). RLHF gathers data with feedback/preference labels, trains a reward model, and then finetunes the LM with reinforcement learning.
@@ 3 Problem Statement
Problem Formulation. In the context of language models, a fine-tuning dataset is a collection of natural language sequences . Each sequence is denoted as , where represents a token. For privacy protection, the users annotate each sequence in the corpus by a binary sequence denoted as , where denotes the token is private tokens (e.g., PII) need to be protected in the context, and otherwise. Here, the contextual privacy posits that the sensitivity of a piece of information is not solely intrinsic to the information itself, but is also influenced by its surrounding context.
To illustrate, “Alan Gates visited Crescent Vale Medical Center for Hemophilia treatment” is considered more indicative than “Alan Gates visited Crescent Vale.” The former provides a clearer insight into an individual’s health when the name “Alan Gates” is paired with the medical condition and the specific medical center. Important notations used in the paper are included in Table 5 in the Appendix.
@@ 4 Methodology
Our methodology adopts a two-pronged approach: 1) corpus curation (i.e. filtering), where sensitive data such as personally identifiable information (PII) is removed from the corpus; and 2) tuning towards the targeted PII-free output. We commence with a theoretical analysis of the information loss incurred by the corpus curation strategy, which provides guidelines for method development. Then, we propose five novel strategies for privacy protection when fine-tuning large language models.
@@ 4.1 Theoretical Analysis on the Information Loss During Corpus Curation
The alternative approach is to mask the text sequence by substituting the word with a special token wherever , then train the model to directly predict the new sequence . Here, denotes a PII token associated with sensitive information like names, organizations, addresses, and website URLs. Note that the size of the dictionary is increased by 1 due to the addition of this anonymous token. The masking procedure above is a one-way mapping from to . We denote this masking mapping as and . The revised maximum log-likelihood estimator is:
@@
This implies that the right-hand side of Eq. 4.1 is larger than the right-hand side of Eq. 4.1. Therefore, directly learning offers richer information. Minimizing Eq. 4.1 ensures the value in Eq. 4.1 remains small, whereas the reverse does not hold. Overall, instructing the model with the “correct” information is more effective and informative than imposing constraints to selectively forget previously acquired knowledge, such as intentionally removing or masking PIIs in the training text.
@@ 4.2.2 Penalty-Based Loss
where and are the penalty terms for generating unigrams and bigrams associated with PII. is the likelihood of generating the token associated with PII at position . is the set of n-grams associated with PII. To construct , we extract all PII-associated n-grams from the training set using scrubadub111https://github.com/LeapBeyond/scrubadub. The cumulative loss is then calculated as:
@@ 4.2.5 DPO
Compared to RLHF, DPO (Rafailov et al., 2023) eliminates the need to train a reward model, and optimizes the same objective as in RLHF with a single stage of policy training using the objective:
@@ 5.1 Datasets
We adopt three biomedical datasets: pii-wikidoc_patient_information, pii-wikidoc, and pii-medical_flashcards as summarized in Table 6 in the Appendix A.1. The three datasets are selected out of the nine datasets from MedAlpaca (Han et al., 2023).
@@ 5.2 Experimental Setup
For each method, we adapt the Alpaca-style tuning pipeline of LLaMA-2 (Touvron et al., 2023), from llama-recipes222https://github.com/facebookresearch/llama-recipes/. In our experiments, all the methods share the same training settings. The number of training epochs is set to 5 and the batch size is 64. For a fair comparison, we adopt the same backbone LLaMA-2 for fine-tuning. More implementation details are included in the Appendix D.